OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4.

OpenStudy (campbell_st):

the general form is \[(x - h)^2 = 4a(y - k)\] the distance from the focus to the directrix is 2a and the vertex is halfway between the directrix and focus on the line of symmetry x = 0 sin the focus is (0, -4) the line of symmetry is x = 0 so 2a = 8 a = 4 the vertex is 4 above the focus as the parabola is concave down graph the information so find the vertex and the equation is \[(x - h)^2 = -4 \times 4 \times (y - k)\] where (h, k) is the vertex. hope it helps